OpenStudy (anonymous):

Why does SO2 and water make H2SO3 and NOT H2SO4

OpenStudy (anonymous):

Well, you can't exactly balance it any other way, the chemical equation is pretty straightforward... SO2+H2O = H2SO3 To get H2SO4, SO2 must be oxidised to form SO3, Sulphur trioxide 2SO2+ O2= 2SO3 SO3 can the dissolve in water to form acid rain: SO3 + H2O = H2SO4